OpenStudy (anonymous):

Simplify log5 s + log5 11 – log5 15 – log5 t

OpenStudy (anonymous):

\[\log_{b} a= \log a/\log b\]

OpenStudy (ash2326):

\[log_{5} s + log_{5} 11 - log_{5} 15 -log_{5} t\] we know that \[log_a(x)+log_a(y)=log_a(xy)\] and \[log_a(x)-log_a(y)=log_a(x/y)\] so using the above two properties \[log_5 s*11-(log_5 15+log_5 t)\] using the addition property for the terms in bracket we get \[log_5s*11-log_5 15*t\] now using the subtraction property we obtain \[log_5 \frac{11*s}{15*t}\]